TO 0 -21 Find bases for the eigenspaces corresponding to the eigenspaces of A = |1 2 1 11 0 3

If ||ul| 4 and v = (2,-1,-2), and if u and v are orthogonal, find ||u + v||.
Answer #1

0 2 3 find Eigen values fint x-s,x2+5)-52=0 b+2+3= 5 S = Trace 52 - a ot 6+ (2) Sy = |A = -2 (0-2) - 4 5+2+87-4 = 0 13 y 1-5= +) 2 -) 3 12 13 2. - = 2,=-2 23 & +213=0 a tha th3 -1, -13 = +223-23 22 xz - 283 X 0 23 az 92 2 2 2 2 1) O 23 1 +23 , -2v u orthogonal 7. 11411 = 4 u v=0 va (2,-1,-2) assume o = 29-b-2c (a, b, c) faltb²tc2 = 4 2 + Ilutull = 2 utu <?+9, -1tb, 2+0